Find the value of n so that parallelogram RSQT is a rectangle if RQ = 2n - 10 and ST = 4n - 30. 20 07 O 10 5
In rectangle LMNP, if LN = 3n + 16 and PM= 5n - 18, find LN. X 67 17 57 05
Answer #1

are two RO and ST diagonals of a rectangle and diagonals of a no are equal Ro = 3T readangle R 2h -10= 4 n-30 30-10=4n- 2h 2LN = = 3n+16 3 x17 +16 67 = LN = 67